LABORATORY AND ASTRONOMICAL DETECTION OF RHOMBOIDAL SiC_{3}

\underline{A. J. APPONI}, M. C. McCARTHY, C. A. GOTTLIEB and P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138.

The rotational spectrum of rhomboidal SiC_{3} in its ^{1}A_{1} electronic ground state was measured in a laboratory discharge of silane and acetylene by Fourier transform microwave spectroscopy and direct absorption millimeter wave spectroscopy. A full structure derived from observations of isotopically substituted species shows that SiC_{3} has a planar geometry with C_{2v} symmetry that consists of atoms in the shape of a distorted four-member ring and a transannular carbon-carbon bond. Following the laboratory detection, seven rotational transitions of SiC_{3} near 100 GHz were detected in the circumstellar envelope of the evolved carbon star IRC+10216 using the NRAO 12 meter telescope.